Find the greatest term in the expansion of (7-5x)^2 where x=2/3
Please explain the steps.
Thankyou in advance.

Grade:12

2 Answers

Nishant Vora IIT Patna
9 years ago

Hello Student. Please find the solution


(7-5x)^2 = 49 + 25x^2 – 70x
so first term = 49
second term = -70*2/3
third term = 25 x^2 = 25*4/9 = 11.11

So 49 is greatest
____________________________________________________________________________________________________________________________________________________________________________________________
Swagat Bordoloi
9 years ago
Sorry sir.. i am extremely sorry.. Actually i typed in the wrong question.. it was (7-5x)^11, where x=2/3.. I am extremely sorry once again. Sir please give me the correct answer for this.